An electrically neutral model airplane is flying in a horizontal circle on a 3.5-m guideline, which is nearly parallel to the ground. The line breaks when the kinetic energy of the plane is 50 J. Reconsider the same situation, except that now there is a point charge of +q on the plane and a point charge of -q at the other end of the guideline. In this case, the line breaks when the kinetic energy of the plane is 53.5 J. Find the magnitude of the charges.
Answer #1

For the centripetal force, we have

Fc = m v2 / r = (2 K.E) / r

so,

Tmax + ke |q|2 / r2 = (2 K.E)charged / r                                           { eq.1 }

And

Tmax = (2 K.E)uncharged / r                                                { eq.2 }

subtracting eq.2 from eq.1 and we get -

ke |q|2 / r2 = 2 [(K.E)charged - (K.E)uncharged] / r

solving for |q|, we have

|q| = 2r [(K.E)charged - (K.E)uncharged] / ke

|q| = 2 (3.5 m) [(53.5 J) - (50 J)] / (9 x 109 Nm2/C2)

|q| = [(24.5 Nm2) / (9 x 109 Nm2/C2)]

|q| = 5.21 x 10-5 C